OpenStudy (anonymous):

The length of one side of a square is 13.5 centimeters. Find the length of a diagonal of the square.

OpenStudy (anonymous):

a^2 + b^2 = c^2 since a triangle is being formed. 13.5^2 + 13.5^c = c^2 c^2 = 364.5 c=\[\sqrt{364.5}\] c=19.09